Can you arrange four 1`s and use of up-to 2 math symbols , make the total be 100 ?

Grade:12

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student

1/.1*1/.1=10*10=100
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty